OpenStudy (anonymous):

simplify (a^3/ab)/(a^5b/a^3b^2)

OpenStudy (anonymous):

\[\frac{a^3}{ab}\times \frac{a^3b^3}{a^5b}\] is a start

OpenStudy (anonymous):

then add up exponents top and bottom, then cancel by subtracting the exponents

OpenStudy (anonymous):

ok let me try that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i had a typo. should have been \[\frac{a^3}{ab}\times \frac{a^3b^2}{a^5b}\]

OpenStudy (anonymous):

i just got b everything else cancels out

OpenStudy (anonymous):

yes in fact if i hadn't made my typo you would have seen that everything cancels and you get 1

OpenStudy (anonymous):

i put a b^3 in the numerator, should have been b^2

OpenStudy (anonymous):

i saw that and i fixed it and got 11 thanks

OpenStudy (anonymous):

hope you got 1, not 11 yw

OpenStudy (anonymous):

i got 1 i hit the button 2x